# How should AI decisions be explained?

### v1

@@
Abstract.
This paper investigates the relationship between law and eXplainable Artificial Intelligence (XAI). While there is much discussion about the AI Act, for which the trilogue of the European Parliament, Council and Commission recently concluded, other areas of law seem underexplored. This paper focuses on European (and in part German) law, although with international concepts and regulations such as fiduciary plausibility checks, the General Data Protection Regulation (GDPR), and product safety and liability. Based on XAI-taxonomies, requirements for XAI-methods are derived from each of the legal bases, resulting in the conclusion that each legal basis requires different XAI properties and that the current state of the art does not fulfill these to full satisfaction, especially regarding the correctness (sometimes called fidelity) and confidence estimates of XAI-methods.
Law, EU, European Law, German Law, eXplainable Artificial Intelligence (XAI), XAI, Interpretable ML, norms, standards, regulation, safe ML
copyright: noneccs: Social and professional topics Governmental regulationsccs: Social and professional topics Management of computing and information systemsccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligence
1. Introduction
The explainability of AI systems is crucial for their development and deployment in sensitive domains. Therefore, the AI Act draft of the European Union (EU) mandates that high-risk AI systems ensure “sufficient transparency” for users to interpret and utilize the system’s results appropriately. However, a challenge lies in the vagueness of these transparency requirements and what exactly constitutes “sufficient transparency”. While more details on this “sufficient transparency” will hopefully be specified in standards such as the ISO/IEC CD TS 6254, other legal bases for the use of eXplainable Artificial Intelligence (XAI) already exist. This paper reviews the legal requirements for explainability of AI systems based on European (and in part German) law, focusing on three legal domains and four application areas. Therefore, our contributions can be summarized as follows:
  • We present an extended list of XAI properties which can be used to formulate requirements.
  • We describe legal bases for the use of XAI and derive requirements such methods need to fulfill in each of the considered areas of law.
@@ 1. Introduction
For this, we first provide a brief overview of common XAI terminology in the next section and establish the choice of areas of law. After the description of related works in Chapter 2, the XAI properties used in this paper are described in Chapter 3, with the distinction of method-specific properties and process-specific properties. In Chapter 4, each of the considered areas of law is introduced, accompanied by the formulation of the task for XAI in this area, from which the XAI requirements are derived. The paper is closed with a discussion in Chapter 5 and a summary in Chapter 6. The paper structure and requirements for XAI are also replicated in short in Table 1 in Chapter 6.
$1.$ Taxonomy
@@
We start by summarizing the taxonomy of explainability methods for AI briefly. For a full introduction we point the readers to (Burkart and Huber, 2021; Molnar, 2022). On the highest level, explanations can be divided into local and global ones. Local explanations explain a particular decision (e.g., the loan was rejected due to a low income) while global explanations reveal how decisions are made in general (e.g., every applicant with income above 100k is given a loan). They are introduced in more detail below. It needs to be noted that most XAI-methods provide non-causal explanations, although special causal methods exist.
1.1.1. Local Explanations
@@
Feature attributions are widely used to provide local explanations by quantifying the contribution of specific features to a decision. For example, in the loan granting scenario, feature attribution could reveal that a short length of employment had the highest negative impact on the rejection, while an income of 75k was the strongest positive feature. Methods like SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) and Linear Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) are commonly used for computing feature attributions. Saliency maps, on the other hand, highlight important regions in images that contribute to specific decisions. They can be created using methods like SHAP or LIME, although usually methods specifically for neural networks—such as Grad-CAM (Selvaraju et al., 2020)—are used. Counterfactual explanations are a type of local explanation that provide insight into decision-making without revealing the detailed functioning of the AI model (Wachter et al., 2018). These explanations show a change in input features necessary to alter the outcome. For instance, a counterfactual explanation for the loan example could be “If you had worked at least five years, you would have been granted the loan.” Data attribution or sample-based explanations involve identifying influential samples from the training set that contributed to a decision. In the loan example, this could be interpreted as “Your loan was rejected because similar applicants from the training data defaulted on a loan,” potentially accompanied by such similar applicants.
1.1.2. Global Explanations
@@
Global explanations provide information about the workings of a model across its entire data range. The simplest form of global explainability is the usage of so-called whitebox models such as decision trees. These models provide out-of-the-box insights into the decision making process and, hence, do not need any specific post-hoc explainability—given that they are small enough in size to be understandable by humans. Providing global explanations for black-box models is more difficult. Apart from partial dependence plots (PDPs), surrogate models provide a way of opening the black box to some degree. For this, either a black-box model is transformed into a white-box one (Nguyen et al., 2020) or a white-box (surrogate) model is fitted on the output of the black-box model that shall be explained, thus showing how similar decisions could be made. As a further way to explain neural networks, especially in computer-vision tasks, concept-based explanations can be used. They usually provide an account of which concepts, e.g., certain patterns like stripes or dots, are especially relevant for the classification of single images (local concept-based explanations) or for the class or model under consideration in general (global concept-based explanations) (Kim et al., 2018). Other methods produce global feature attribution explanations (Covert et al., 2020), highlighting features with high predictive power.
1.2. Considered Laws
This paper focuses on EU and German laws, although with a wide span of possible stakeholders and reasons for the use of XAI (see Chapter 4). While there is uncertainty about what XAI can achieve, the considered areas of law—company law, the General Data Protection Regulation (GDPR), product safety and liability for defect products—are among the most discussed and show the different legal questions XAI could possibly answer. Additionally, these areas bear similarity across many different jurisdictions, as the ISION-rule (as used in Chapter 4.1) serves as a specification of fiduciary duties, which are applicable to the directors and officers of nearly every corporation worldwide. Moreover, the ISION-rule is closely connected to § 141(e) of the Delaware General Corporation Law (Title 8, Chapter 1 of the Delaware Code) in the US, addressing the reliance on expert opinions by board directors. The GDPR (Chapter 4.2) was introduced as a de-facto international standard, coining the term “Brussels-Effect” for EU laws that are adhered to even outside the EU. Product safety and liability (Chapter 4.3) are harmonized within the EU. This paper does not intend to provide an exhaustive list of legal bases for the use of XAI, but aims to provide a starting point towards such a list and shows how laws can inform the development of XAI techniques.
2. Related Work
@@
To date, not many works have concerned themselves with the link between law and XAI, apart from works around the AI Act and the GDPR. In the following, the related works and the understanding of transparency in law are summarized. A broader discussion around the need for algorithmic transparency was started by the provisions of the GDPR, especially about whether a “right to explanation” exists (Wachter et al., 2017) and how much information would be necessary to fulfill such a right. Wachter et al. (2018) devised counterfactual explanations to give users actionable advise while keeping the exact workings of an Machine Learning (ML) model a secret (see Section 1.1.1). Vorras and Mitrou (2021) argue for a “right to explanation” with the implication of a “qualified transparency mechanism”, i.e., adapting explanations to the target audience and “considering the technical limitations of AI interpretability, the underlying legal and regulatory barriers, as well as the legitimate expectations of stakeholders”. A similar sentiment can be found in (Błażej Kuźniacki, Marco Almada, Kamil Tyliński, Łukasz Górski, Beata Winogradska, Reza Zeldenrust, 2022), where Kuzniacki et al. argue for the construction of a global regulatory framework for the use of XAI in tax law. They additionally recommend that “tax lawyers and computer scientists should work together to help reach a consensus on legal and technical solutions that secure XAI’s place in tax law”. While our work is not based on tax law, it follows a similar idea as it aims to combine the information from legal experts and computer scientists to make progress towards a structured use of XAI for legal purposes. As of now, a gap between the legal and technical understanding can be discerned in many works, as echoed by Gyevnar et al. (2023), who argue that this difference in understanding can also be seen in the EU AI Act. They note that there seem to be multiple understandings of transparency and that the transparency mandated in the EU AI Act is different from purely technical transparency, which is the aim of most works in XAI. Also concerned with the different interpretations of transparency in the AI Act are Madalina Busuioc, Deirdre Curtin, Marco Almada (2022). They stress that the explainability of AI systems, i.e., the presentation of specific—mediated—details of their workings, is different from the disclosure of the model itself, where only the second option allows for the unbiased scrutiny of such systems. Such explanations and thus, mediated information, could be used maliciously, as shown by Zhou and Joachims (2023). Summarizing, despite some works in that direction, the use of XAI for legal purposes is still unclear.
3. Properties of XAI-Methods (possibly) relevant for their legal use

### v2

@@
Abstract
This paper investigates the relationship between law and eXplainable Artificial Intelligence (XAI). While there is much discussion about the AI Act, which was adopted by the European Parliament in March 2024, other areas of law seem underexplored. This paper focuses on European (and in part German) law, although with international concepts and regulations such as fiduciary duties, the General Data Protection Regulation (GDPR), and product safety and liability. Based on XAI-taxonomies, requirements for XAI methods are derived from each of the legal fields, resulting in the conclusion that each legal field requires different XAI properties and that the current state of the art does not fulfill these to full satisfaction, especially regarding the correctness (sometimes called fidelity) and confidence estimates of XAI methods.
1 Introduction
The AI Act of the European Union (EU) mandates that the design of high-risk AI systems enables deployers to “interpret a system’s output” (Art. 13(1) and 13(3)b vii of the AI Act) and notes that, for example in the case of AI use in law enforcement the “exercise of important procedural fundamental rights []delimited-[][...][ … ] could be hampered, in particular, where such AI systems are not sufficiently transparent, explainable and documented“ (Recital 59 of the AI Act). However, a challenge lies in the vagueness of these transparency requirements and what exactly constitutes “sufficiently” transparent and explainable. While more details on this will hopefully be specified in standards such as the ISO/IEC CD TS 6254, other legal bases for the use of eXplainable Artificial Intelligence (XAI) already exist. This paper reviews the legal requirements for explainability of AI systems based on European (and in part German) law, focusing on three legal domains and application areas. Therefore, our contributions can be summarized as follows:
  • We present an extended list of XAI properties that can be used to formulate requirements.
  • We describe legal bases for the use of XAI—besides the well-known case of the GDPR, also fiduciary duties and product safety and liability—and derive requirements that XAI methods need to fulfill in each of the considered areas of law.
@@ 1 Introduction
For this, we first provide a brief overview of common XAI terminology in the next section and establish the choice of areas of law. After the description of related works in Section 2, the XAI properties used in this paper are described in Section 3, with the distinction of method-specific properties and process-specific properties. Based on the separation of decision-centric and model-centric approaches, Section 4 and Section 5 describe the relevant legal fields and formulate a task for XAI in each of the relevant application areas. Based on this task, the requirements for XAI methods are derived. The paper closes with a discussion in Section 6 and a summary in Section 7. The paper structure and requirements for XAI are also replicated in short in Table 2 in Section 7.
1.1 Taxonomy
@@
We start by summarizing the taxonomy of explainability methods for AI briefly. For a full introduction we point the readers to (Burkart and Huber 2021; Molnar 2022). At the highest level, explanations can be divided into local and global ones. Local explanations explain a particular decision (e.g., the loan was rejected due to a low income) while global explanations reveal how decisions are made in general (e.g., every applicant with an income above €100k is given a loan). They are introduced in more detail below. It needs to be noted that most XAI methods provide non-causal explanations, although special causal methods exist.
Local Explanations
@@
Feature attributions are widely used to provide local explanations by quantifying the contribution of specific features to a decision. For example, in the loan granting scenario, feature attribution could reveal that a short length of employment had the highest negative impact on the rejection, while an income of €75k was the strongest positive feature. Methods like SHapley Additive exPlanations (SHAP) (Lundberg and Lee 2017) and Linear Interpretable Model-Agnostic Explanations (LIME) (Ribeiro, Singh, and Guestrin 2016) are commonly used for computing feature attributions. Saliency maps, on the other hand, highlight important regions in images that contribute to specific decisions. They can be created using methods like SHAP or LIME, although usually methods specifically for neural networks—such as Grad-CAM (Selvaraju et al. 2020)—are used. Counterfactual explanations are a type of local explanation that provide insight into decision-making without revealing the detailed functioning of the Machine Learning (ML) model (Wachter, Mittelstadt, and Russell 2018). These explanations show a change in input features necessary to alter the outcome. For instance, a counterfactual explanation for the loan example could be “If you had worked at least five years, you would have been granted the loan.” Data attribution or sample-based explanations involve identifying influential samples from the training set that contributed to a decision. In the loan example, this could be interpreted as “Your loan was rejected because similar applicants from the training data defaulted on a loan,” potentially accompanied by such similar applicants.
Global Explanations
@@
Global explanations provide information about the workings of a model across its entire data range. The simplest form of global explainability is the usage of so-called white-box models such as decision trees. These models provide out-of-the-box insights into the decision making process and, hence, do not need any specific post-hoc explainability—given that they are small enough in size to be comprehensible to humans. Providing global explanations for black-box models is more difficult. Apart from partial dependence plots (PDPs), surrogate models provide a way of opening the black box to some degree. For this, either a black-box model is transformed into a white-box one (Nguyen, Kasmarik, and Abbass 2021) or a white-box (surrogate) model is fitted on the output of the black-box model that shall be explained, thus showing how similar decisions could be made. As a further way to explain neural networks, especially in computer-vision tasks, concept-based explanations can be used. They usually provide an account of which concepts, e.g., certain patterns like stripes or dots, are especially relevant for the classification of single images (local concept-based explanations) or for the class or model under consideration in general (global concept-based explanations) (Kim et al. 2018). Other methods produce global feature attribution explanations (Covert, Lundberg, and Lee 2020), highlighting features with high predictive power.
1.2 Legal Necessity for the Use of XAI
To date, most works about the link between law and XAI focus on the GDPR or the AI Act. In reality, the relationship between law and XAI is even closer. There are many situations in which the law requires explanations. In this paper, we focus on the need for explanation in relations between private actors, although such a need also exists in relations between the citizen and the state (for example, in the adoption of automated administrative acts or in criminal justice). In doing so, we will distinguish between the need for global and for local explanations in the sense of the taxonomy outlined above. Whether global or local explanations are required depends on the rights, obligations and legally protected interests involved. If, for example, a loan application is rejected based on the ML calculation, the applicant will want to know the reasons for this in their specific case. If a “right to explanation” really exists in this situation (e.g., under the GDPR), there is a legal need for (at least) a local explanation. If, on the other hand, the safety of an ML application is at stake, a global explanation may be desirable from a legal perspective. More generally, it depends on whether the legal focus is on a single decision or the entire model. For each of these two groups of cases, we will present representative subgroups or examples which, by their very nature, do not claim to be exhaustive. The need for explainability of an individual decision is considered (1) from the perspective of the decision-makers, whereby we selected company directors as a representative group (company law) and (2) from the perspective of the decision recipients (data protection law). When discussing the need for explainability of the entire model, the focus will be on (3) developers (product safety law) and (4) operators (product liability). This choice can be mapped to the definition of stakeholders for whom explanations are created described by Tomsett et al. (2018) as operators and executors (group 1), data-subjects and decision-subjects (group 2), and creators and examiners of AI systems (group 3 and 4). Note that an often discussed use case for XAI is criminal litigation, which is not explicitly described here, because the requirements there depend on the specific question asked. Thus, for criminal litigation, all explanation types could be used, given the right litigation task. As an example, in one of the first cases of XAI-use in litigation, feature importance explanations with varying degrees of access to model and data were used (Fraser, Simcock, and Snoswell 2022).
2 Related Work
As described before, most related works have concerned themselves with the link between law and XAI regarding the AI Act and the GDPR. In the following, some of these works and the understanding of transparency in law are summarized.
@@ 2 Related Work
A broader discussion around the need for algorithmic transparency was started by the provisions of the GDPR, especially about whether a “right to explanation” exists (Wachter, Mittelstadt, and Floridi 2017) and how much information would be necessary to fulfill such a right. Wachter, Mittelstadt, and Russell (2018) devised counterfactual explanations to give users actionable advise while keeping the exact workings of an ML model a secret (see Section 1.1). Vorras and Mitrou (2021) argue for a “right to explanation” with the implication of a “qualified transparency mechanism”, i.e., adapting explanations to the target audience and “considering the technical limitations of AI interpretability, the underlying legal and regulatory barriers, as well as the legitimate expectations of stakeholders”. A similar sentiment can be found in (Błażej Kuźniacki, Marco Almada, Kamil Tyliński, Łukasz Górski, Beata Winogradska, Reza Zeldenrust 2022), where Kuzniacki et al. argue for the construction of a global regulatory framework for the use of XAI in tax law. They additionally recommend that “tax lawyers and computer scientists should work together to help reach a consensus on legal and technical solutions that secure XAI’s place in tax law”. While our work is not based on tax law, it follows a similar idea as it aims to combine the information from legal experts and computer scientists to make progress towards a structured use of XAI for legal purposes. As of now, a gap between the legal and technical understanding can be discerned in many works, as echoed by Gyevnar, Ferguson, and Schafer (2023), who argue that this difference in understanding can also be seen in the EU AI Act. They note that there seem to be multiple understandings of transparency and that the transparency mandated in the EU AI Act is different from purely technical transparency, which is the aim of most works in XAI. Also concerned with the different interpretations of transparency in the AI Act are Madalina Busuioc, Deirdre Curtin, Marco Almada (2022). They stress that the explainability of AI systems, i.e., the presentation of specific—mediated—details of their workings, is different from the disclosure of the model itself, where only the second option allows for proper scrutiny of such systems. Such explanations and thus, mediated information, could be used maliciously, as shown by Zhou and Joachims (2023). Summarizing, despite some works in that direction, the use of XAI for legal purposes is still unclear.
3 Properties of XAI-Methods (Possibly) Relevant for their Legal Use